1a) The standard cost card of a particular product specifies that it requires 3.0 direct labor-hours...

1a)

The standard cost card of a particular product specifies that it requires 3.0 direct labor-hours at $10.20 per direct labor-hour. During March, 4,000 units of the product were produced and direct labor wages of $120,750 were incurred. A total of 15,000 direct labor-hours were worked. The direct labor variances for the month were:

Labor Rate Variance Labor Efficiency Variance
A) $1,650 F $28,124 U
B) $1,650 F $30,600 U
C) $32,250 F $28,124 U
D) $32,250 F $30,600 U

Option A

Option B

Option C

Option D

1b)

Fruchter Corporation keeps careful track of the time required to fill orders. The times recorded for a particular order appear below:

Hours
Move time 26.7
Wait time 1.9
Queue time 0.3
Process time 3.9
Inspection time 10.9


The throughput time was:

rev: 08_01_2016_QC_CS-56553

37.6 hours

43.7 hours

41.8 hours

6.1 hours
Answer #1
Labor Rate Variance = 120750-(15000*10.20)= $32,250 F
Labor Efficiency Variance = 10.20*(15000-4000*3)= $30,600 U
Option D $32,250 F and $30,600 U is correct
1b
Throughput time = Move time +Queue time +Process time +Inspection time
Throughput time = 26.7+0.3+3.9+10.9= 41.8 hours
Option 3 41.8 hours is correct